# Temporal Wasserstein Imputation: Versatile Missing Data Imputation for Time Series

### v1

@@ 3.2 Alternating minimization algorithm
The optimization landscapes of (4) and (5) are complicated because the optimal transport cost is itself the minimum value of a linear program (3), and there are potentially many variables to optimize. This makes directly optimizing (4) or (5) computationally intractable. To remedy this issue, we propose to solve the nested minimization problem using alternating minimization. Since (4) is a special case of (5) with $\lambda=0$, we focus our discussion on (5).
To begin, observe that with an initialized $\hat{\mathbf{w}}^{(0)}$, problem (3) is exactly the discrete optimal transport problem. The minimizer $\hat{\mathbf{\Pi}}$ can be efficiently computed using existing algorithms (Cuturi,, 2013; Peyré and Cuturi,, 2019). Thus, at iteration $k=1,2,\ldots$, treating $\mathbf{w}=\hat{\mathbf{w}}^{(k-1)}$ as fixed, we may compute the optimizer $\hat{\mathbf{\Pi}}^{(k)}$ of (3), which corresponds to the inner optimization problem of (5). Next, we optimize (5) over $\mathbf{w}$, taking $\mathbf{\Pi}=\hat{\mathbf{\Pi}}^{(k)}$ as fixed. This procedure is repeated until convergence. Let
@@ 4      Break if convergence criterion is met and set $\hat{\mathbf{w}}=\hat{\mathbf{w}}^{(k)}$
Next, we investigate the properties of Algorithm 1 in an important special case where closed forms can be obtained. Suppose $L(\mathbf{v}_{i},\mathbf{v}_{j})=\|\mathbf{v}_{i}-\mathbf{v}_{j}\|^{2}$. Then it is not difficult to show that $F(\mathbf{w},\mathbf{\Pi})=\mathbf{w}^{\top}\mathbf{H}(\mathbf{\Pi})\mathbf{w}$, where
$\mathbf{H}(\mathbf{\Pi})$ $=\sum_{h=0}^{p-1}\mathbf{A}_{h}(\mathbf{\Pi})+\frac{\lambda}{2}\mathbf{I}_{n},\mbox{ and}$
@@
in which all unspecified entries in $\mathbf{A}_{h}(\mathbf{\Pi})$ are zero, and $\mathbf{0}_{h}$ is an $h\times h$ matrix of zeros. Hence $F$ is quadratic in $\mathbf{w}$. In fact, it can be shown that $\mathbf{H}(\mathbf{\Pi})$ is nonnegative definite for any $\mathbf{\Pi}$ satisfying the coupling constraints. Moreover, $F$ is strongly convex in $\mathbf{w}$ if $\lambda>0$ (see also Proposition 4 in Section 4). Thus many efficient solvers are applicable in this step (Boyd and Vandenberghe,, 2004; Parikh et al.,, 2014). If, in addition, the constraint set $\mathcal{C}$ is defined through a system of linear equations, i.e.,
$\mathcal{C}=\{\mathbf{w}\in\mathbb{R}^{n}:\mathbf{Kw}=\mathbf{b}\}$
for some known $\mathbf{K}$ and $\mathbf{b}$, then step (b) in Algorithm 1 has the following closed form
@@
In contrast, (8) shows that the temporal Wasserstein imputation (with $p=2$) is given by
@@
which is a convex combination of all observed values. This is somewhat expected because, as a nonparametric method, TWI shall use all data to learn the underlying dynamics.
@@
$\hat{\mathbf{w}}\in$ $\arg\min_{\mathbf{w}\in\mathcal{C}}F(\mathbf{w},\hat{\mathbf{\Pi}}),$
$\hat{\mathbf{\Pi}}\in$ $\arg\min_{\mathbf{\Pi}}F(\hat{\mathbf{w}},\mathbf{\Pi}).$
In addition, suppose $\lambda=0$ for a moment. Let $\mathbf{u}=(u_{0},u_{1},\ldots,u_{p-1})$ be a random vector whose distribution is $\hat{\mu}_{\mathrm{pre}}$ and $\mathbf{v}=(v_{0},\ldots,v_{p-1})$ a random vector whose distribution is $\hat{\mu}_{\mathrm{post}}$. Furthermore, let the joint distribution of $(\mathbf{u},\mathbf{v})$ be given by the optimal transport coupling between $\hat{\mu}_{\mathrm{pre}}$ and $\hat{\mu}_{\mathrm{post}}$, which is $\hat{\mathbf{\Pi}}$ (by viewing it as a joint distribution between the two discrete measures). Consider an index $s\in\mathcal{M}$ with $p-1\leq s\leq n_{1}-p+1$. By (8), it is not difficult to see that
@@
$=$ $\frac{1}{p}\sum_{i=s}^{s+p-1}\sum_{j=n_{1}+1}^{n-1}\mathbb{P}_{\hat{\mathbf{\Pi}}}\left(\mathbf{v}=(\hat{w}_{j},\ldots,\hat{w}_{j-p+1})|\mathbf{u}=(\hat{w}_{i},\ldots,\hat{w}_{i-p+1})\right)\hat{w}_{j-i+s}$
$=$ $\frac{1}{p}\sum_{h=0}^{p-1}\mathbb{E}_{\hat{\mathbf{\Pi}}}(v_{h}|\mathbf{u}=(\hat{w}_{s+h},\ldots,\underbrace{\hat{w}_{s}}_{\mbox{\makebox[0.0pt]{$h$-th element}}},\ldots,\hat{w}_{s+h-p+1})).$ (9)
@@
Remark 3.
(Barycenter interpretation) The summands in (9) is also known as the barycentric projections (Ambrosio et al.,, 2006; see also Deb et al.,, 2021). Hence the temporal Wasserstein imputation can also be interpreted as an average of barycentric projections under the coupling $\hat{\mathbf{\Pi}}$.
Because of the optimization formulation, it is straightforward to extend TWI to multivariate time series. Suppose the multivariate time series of interest is $\mathbf{x}_{t}=(x_{t,1},\ldots,x_{t,d})^{\top},t=0,1,\ldots,n-1$. To compute the imputation $\mathbf{W}=(\mathbf{w}_{0},\mathbf{w}_{1},\ldots,\mathbf{w}_{n-1})$ where $\mathbf{w}_{t}=(w_{t,1},\ldots,w_{t,d})^{\top}$, we may replace $\mathbf{v}_{i}(\mathbf{w})$ in (3) by
$\mathbf{v}_{i}(\mathbf{W}):=(w_{i,1},\ldots,w_{i-p+1,1},w_{i,2},\ldots,w_{i-p+1,2},\ldots,w_{i-p+1,d})^{\top}.$
Then Algorithm 1 can still be applied. Since the multivariate information is summarized by the distance matrix $\mathbf{L}=(l_{ij})_{0\leq i\leq n_{1},n_{1}+1\leq j\leq n-1}$, where $l_{ij}=L(\mathbf{v}_{i}(\mathbf{W}),\mathbf{v}_{j}(\mathbf{W}))$, regardless of the dimension $d$, step (a) in Algorithm 1 is identical as the scalar case. In the notable special case where $L(\mathbf{u},\mathbf{v})=\|\mathbf{u}-\mathbf{v}\|^{2}$, (6) becomes
@@
where $\|\cdot\|_{F}$ is the Frobenius norm. Straightforward calculations show that fixing $\mathbf{\Pi}=\hat{\mathbf{\Pi}}$, if $\hat{\mathbf{W}}$ is a solution to step (b) of Algorithm 1, then for each $l\in\{1,2,\ldots,d\}$, $(\hat{w}_{0,l},\hat{w}_{1,l},\ldots,\hat{w}_{n-1,l})$ also satisfies the system of linear equations (8) in the case of no side information. Therefore, in step (b) of Algorithm 1, each component in the multivariate time series can be imputed in parallel, and the imputation only depends on the future and past values of individual series. The cross-sectional information is only exploited when learning the optimal transport coupling in step (a). Thus, the algorithm is fast to implement in the multivariate setting.
We shall compare TWI with the approach proposed by Carrizosa et al., (2013), which also aims to alleviate the corrupted statistical patterns caused by existing imputation methods. Their method corrects the resulting statistical pattern by penalizing the deviation of some moments from the theoretical values. For instance, let $\rho_{k}$, $k=1,2,\ldots,k_{0}$ be some pre-specified values for the autocorrelation function at lag $k$, and $m_{k}$, $k=1,2,\ldots,k_{1}$, the pre-specified values for the $k$-th moments. Then imputation is done by solving the following program.
$\min_{\mathbf{w}\in\mathcal{C}}\frac{f(\mathbf{w})}{f(\mathbf{w}^{(0)})}+\sum_{k=1}^{k_{0}}\lambda_{k}\left(\frac{\rho_{k}(\mathbf{w})}{\rho_{k}}-1\right)^{2}+\sum_{k=1}^{k_{1}}\mu_{k}\left(\frac{n^{-1}\sum_{t=0}^{n-1}w_{t}^{k}}{m_{k}}-1\right)^{2},$
@@
where $\rho_{k}(\mathbf{w})$ is the sample autocorrelation function computed from the series $\mathbf{w}$, $f(\mathbf{w})=\sum_{t=1}^{n-1}(w_{t}-w_{t-1})^{2}$, and $\mu_{k},\lambda_{k}\geq 0$ are the penalty strengths. As a result, the moments and the autocorrelation functions of the imputed series will be close to the pre-specified values. However, this approach has a number of drawbacks. First, a good theoretical value for $\rho_{k}$ and $m_{k}$ may not be available in practice, especially when the missing values are many. Second, minimizing $f$ promotes smoothness and a positive correlation between $w_{t}$ and $w_{t-1}$, which does not necessarily align with the underlying time series. Finally, the optimization problem is highly non-convex, leading to intense computational hurdle. In contrast, the temporal Wasserstein imputation is data-driven and completely nonparametric, requiring small amount of tuning. It also enjoys the computational advantages brought by the alternating minimization algorithm and recent advances in computational optimal transport.
@@
In closing this subsection, we discuss the imputation of nonstationary time series. For an $I(1)$ process $\{x_{t}\}$, $y_{t}=x_{t}-x_{t-1}$ is stationary and we can apply imputation methods to the first-differenced series $\{y_{t}\}$. In this case, the number of missing data multiplies in the first-differenced series and could compromise the imputation performance. However, for TWI we can mitigate this issue by encoding the information from the raw data via the admissible set $\mathcal{C}$. For example, if $x_{t}$ is observed, then $y_{1}+y_{2}+\ldots+y_{t}=x_{t}-x_{0}$, so we may restrict the imputation $\{w_{t}\}$ for $\{y_{t}\}$ to satisfy $w_{1}+\ldots+w_{t}=x_{t}-x_{0}$ (assuming $x_{0}$ is observed). Finally, although we have motivated TWI from the stationarity of the underlying process, we found that it fares quite well with certain nonstationary processes such as cyclic time series. In Section 5, we further explore the application of TWI to nonstationary time series via simulation studies.
@@ 4 Theoretical properties
In this section, we investigate the optimization and statistical properties of TWI. Under some general assumptions, we first show Algorithm 1 converges to critical points. Then, the asymptotic consistency of TWI is established when the missing data concentrate on one side of $n_{1}$. For general missing patterns, we show that identification of the underlying marginal distributions is possible under the assumptions of asymmetric missing pattern and interpolator stability (see (C1) and (C2) in Section 4.2) using a two-state Markov chain example.
In the following, we maintain the assumption that the index set $\mathcal{M}=\mathcal{M}_{n}$, which corresponds to the missing data, is fixed and independent of the underlying time series. It is worth noting that we allow the number of missing entries to scale with the sample size, i.e., $\sharp(\mathcal{M}_{n})\rightarrow\infty$ as $n\rightarrow\infty$, where $\sharp(E)$ is the cardinality of a set $E$. This asymptotic regime is more suitable for applications where non-negligible portions of the data are missing. In addition, our focus is on the distributional bias induced by the imputations. If $\sharp(\mathcal{M}_{n})=o(n)$, the problem becomes trivial because the empirical marginal distributions of any (bounded) imputations satisfy
@@
where $\mathcal{A}_{n}=\{t:\{t,t-1,\ldots,t-p+1\}\cap\mathcal{M}_{n}\neq\emptyset,t\in\{0,1,\ldots,n-1\}\}$.
@@ 4.1 Optimization and statistical guarantees
We begin by studying the optimization properties of Algorithm 1.
@@
Suppose $\mathcal{C}$ is a convex set. Then,
@@
    (monotonicity) $F(\mathbf{w}^{(t)},\mathbf{\Pi}^{(t+1)})\leq F(\mathbf{w}^{(t)},\mathbf{\Pi}^{(t)})\leq F(\mathbf{w}^{(t-1)},\mathbf{\Pi}^{(t)})$, $t=1,2,\ldots$;
@@
    every limit point $(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})$ of $\{(\mathbf{w}^{(t+1)},\mathbf{\Pi}^{(t)}):t=1,2,\ldots\}$ is a critical point of $F$. That is,
@@
Proof.
Moreover,
$\frac{\partial^{2}F_{0}}{\partial\mathbf{u}^{2}}=$ $\mathrm{diag}\left(\sum_{j=n_{1}+1}^{n-1}\pi_{ij}\ell^{\prime\prime}(w_{i}-w_{j}):i=p-1,\ldots,n_{1}\right),$ (10)
$\frac{\partial^{2}F_{0}}{\partial\mathbf{v}^{2}}=$ $\mathrm{diag}\left(\sum_{i=p-1}^{n_{1}}\pi_{ij}\ell^{\prime\prime}(w_{i}-w_{j}):j=n_{1}+1,\ldots,n-1\right),$ (11)
$\frac{\partial^{2}F_{0}}{\partial\mathbf{u}\partial\mathbf{v}}=$ $-\left(\pi_{ij}\ell^{\prime\prime}(w_{i}-w_{j})\right)_{p-1\leq i\leq n_{1},n_{1}+1\leq j\leq n-1}.$ (12)
Since $\ell$ is convex, ′′$\ell^{\prime\prime}\geq 0$. From (10)–(12), it is readily seen that the Hessian matrix $\mathbf{H}=\partial^{2}F_{0}/\partial\mathbf{w}^{2}$ is weakly diagonally dominant with nonnegative diagonal entries. It follows from the Gershgorin circle theorem that $\mathbf{H}$ is positive semi-definite. Thus $F_{0}$ is convex in $\mathbf{w}$. ∎
Remark 5.
It is straightforward to generalize Proposition 4 for multivariate time series. Details are omitted.
Due to the biconvexity of $F$, one can employ many off-the-shelf solvers for the subproblems in Algorithm 1. Moreover, it follows directly from the proof that if $\lambda>0$, $F(\mathbf{w},\mathbf{\Pi})$ is $\lambda$-strongly convex in $\mathbf{w}$. Hence, step (b) in Algorithm 1 admits a unique solution in each iteration and can be solved efficiently.
In general, not all critical points of $F$ yield desirable imputations, and even its global minimizer may not correspond to the correct marginal distributions. For (an extreme) example, imputing a fixed constant when all data are missing results in a zero transport cost, which is clearly arbitrary. However, when the missing data happen to lie on the same side of $n_{1}$, the global minimizer of $F$ achieves asymptotic consistency in the sense of Proposition 6 below. Before stating the proposition, we make the following assumptions, which are satisfied by many linear and nonlinear time series models.
The underlying time series $\{\tilde{x}_{t}\}$ is stationary and ergodic.
$\mathbb{E}|\tilde{x}_{t}|^{k}<\infty$ for some $k\geq 1$.
@@
Assume (A1) and (A2) hold. Suppose for any $\mathbf{w}\in\mathcal{C}$, $w_{t}=\tilde{x}_{t}$ if $t\notin\mathcal{M}_{n}$. Suppose further $n$ and $n_{1}$ satisfy $n_{1}\rightarrow\infty$ and $n-n_{1}\rightarrow\infty$. If either $\mathcal{M}_{n}\subseteq\{0,1,\ldots,n_{1}\}$ or $\mathcal{M}_{n}\subseteq\{n_{1}+1,\ldots,n-1\}$ (but not both), then, the global minimizer $(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})$ of $F$, subject to relevant constraints such as the coupling conditions and with $\lambda=0$ and $L(\mathbf{u},\mathbf{v})=\|\mathbf{u}-\mathbf{v}\|^{k}$, satisfies
$\mathcal{W}_{k}\left(\frac{1}{n_{1}-p+2}\sum_{t=p-1}^{n_{1}}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})},\mu_{p}\right)+\mathcal{W}_{k}\left(\frac{1}{n-n_{1}-1}\sum_{t=n_{1}+1}^{n-1}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})},\mu_{p}\right)\rightarrow 0\quad\mbox{almost surely},$
where $\mu_{p}$ is the $p$-dimensional marginal distribution of $\{\tilde{x}_{t}\}$.
Proposition 6 implies the empirical marginal distributions of the imputed time series converge to the underlying marginal distribution in the Wasserstein distance. This implies
$\hat{\mu}_{\mathrm{pre}}(\hat{\mathbf{w}})=\frac{1}{n_{1}-p+2}\sum_{t=p-1}^{n_{1}}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})}\Rightarrow\mu_{p},\mbox{ and}\quad\hat{\mu}_{\mathrm{post}}(\hat{\mathbf{w}})=\frac{1}{n-n_{1}-1}\sum_{t=n_{1}+1}^{n-1}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})}\Rightarrow\mu_{p}.$
In fact, convergence in Wasserstein distance further ensures for any continuous function $\varphi$ with $|\varphi(x)|\leq C(1+|x|^{k})$ for all $x\in\mathbb{R}$ and some $C<\infty$, we have
@@
[…]
In practice, we may not have all missing data lying on the same side of $n_{1}$, even though we have some freedom to choose $n_{1}$. Unfortunately, asymptotic consistency results for general missing pattern that are in similar spirits as Proposition 6 can hardly be derived. The reason, as we shall see in the next subsection, is that matching the underlying marginal distribution by equating the empirical marginal distributions essentially suffers from an identification problem.

### v2

@@ 3.2 Alternating minimization algorithm
The optimization landscapes of (3) and (4) are complicated because the optimal transport cost is itself the minimum value of a linear program (2), and there are potentially many variables to optimize. To tackle this issue, we propose to solve the nested minimization problem using alternating minimization. Since (3) is a special case of (4) with $\lambda=0$, we focus our discussion on (4).
To begin, observe that with an initialized $\hat{\mathbf{w}}^{(0)}$, problem (2) is exactly the discrete optimal transport problem. Thus, at iteration $k=1,2,\ldots$, treating $\mathbf{w}=\hat{\mathbf{w}}^{(k-1)}$ as fixed, we may compute the optimizer $\hat{\mathbf{\Pi}}^{(k)}$ of (2) efficiently by some existing algorithms (Cuturi,, 2013; Peyré and Cuturi,, 2019). This corresponds to the inner optimization problem of (4). Next, we optimize (4) over $\mathbf{w}$, taking $\mathbf{\Pi}=\hat{\mathbf{\Pi}}^{(k)}$ as fixed. This procedure is repeated until convergence. Let
@@ 4      Break if convergence criterion is met and set $\hat{\mathbf{w}}=\hat{\mathbf{w}}^{(k)}$
Next, we investigate the properties of Algorithm 1 in an important special case where closed forms can be obtained. Suppose $L(\mathbf{v}_{i},\mathbf{v}_{j})=\|\mathbf{v}_{i}-\mathbf{v}_{j}\|^{2}$. Then it is not difficult to show that $F(\mathbf{w},\mathbf{\Pi})=\mathbf{w}^{\top}\mathbf{H}(\mathbf{\Pi})\mathbf{w}$, where $\mathbf{H}(\mathbf{\Pi})=\sum_{h=0}^{p-1}\mathbf{A}_{h}(\mathbf{\Pi})+\frac{\lambda}{2}\mathbf{I}_{n}$ and
@@
in which all unspecified entries in $\mathbf{A}_{h}(\mathbf{\Pi})$ are zero, and $\mathbf{0}_{h}$ is an $h\times h$ matrix of zeros. Hence $F$ is quadratic in $\mathbf{w}$. In fact, it can be shown that $\mathbf{H}(\mathbf{\Pi})$ is nonnegative definite for any $\mathbf{\Pi}$ satisfying the coupling constraints. Moreover, $F$ is strongly convex in $\mathbf{w}$ if $\lambda>0$ (see also Proposition 3 in Section 4). Thus many efficient solvers are applicable in this step (Boyd and Vandenberghe,, 2004; Parikh et al.,, 2014). If, in addition, the constraint set $\mathcal{C}$ is defined through a system of linear equations, i.e., $\mathcal{C}=\{\mathbf{w}\in\mathbb{R}^{n}:\mathbf{Kw}=\mathbf{b}\}$ for some known $\mathbf{K}$ and $\mathbf{b}$, then step (b) in Algorithm 1 has the following closed form
@@
In contrast, (7) shows that the temporal Wasserstein imputation (with $p=2$ and $\lambda=0$) is given by
@@
which is a convex combination of all observed values. This is somewhat expected because, as a nonparametric method, TWI uses all data to learn the underlying dynamics.
@@
$\hat{\mathbf{w}}\in\arg\min_{\mathbf{w}\in\mathcal{C}}F(\mathbf{w},\hat{\mathbf{\Pi}}),\quad\hat{\mathbf{\Pi}}\in\arg\min_{\mathbf{\Pi}}F(\hat{\mathbf{w}},\mathbf{\Pi}).$
In addition, suppose $\lambda=0$ for a moment. Let $\mathbf{u}=(u_{0},u_{1},\ldots,u_{p-1})$ be a random vector whose distribution is $\hat{\mu}_{\mathrm{pre}}$ and $\mathbf{v}=(v_{0},\ldots,v_{p-1})$ a random vector whose distribution is $\hat{\mu}_{\mathrm{post}}$. Furthermore, let the joint distribution of $(\mathbf{u},\mathbf{v})$ be given by the optimal transport coupling between $\hat{\mu}_{\mathrm{pre}}$ and $\hat{\mu}_{\mathrm{post}}$, identified as $\hat{\mathbf{\Pi}}$ (by viewing it as a joint distribution between the two discrete measures). Consider an index $s\in\mathcal{M}$ with $p-1\leq s\leq n_{1}-p+1$. By (7), it is not difficult to see that
@@
$=$ $\frac{1}{p}\sum_{h=0}^{p-1}\mathbb{E}_{\hat{\mathbf{\Pi}}}(v_{h}|\mathbf{u}=(\hat{w}_{s+h},\ldots,\hat{w}_{s},\ldots,\hat{w}_{s+h-p+1})).$ (8)
@@
Because of the optimization formulation, it is straightforward to extend TWI to multivariate time series. Suppose the multivariate time series of interest is $\mathbf{x}_{t}=(x_{t,1},\ldots,x_{t,d})^{\top}$. To compute the imputation $\mathbf{W}=(\mathbf{w}_{0},\mathbf{w}_{1},\ldots,\mathbf{w}_{n-1})$ where $\mathbf{w}_{t}=(w_{t,1},\ldots,w_{t,d})^{\top}$, we may replace $\mathbf{v}_{i}(\mathbf{w})$ in (2) by $\mathbf{v}_{i}(\mathbf{W}):=(w_{i,1},\ldots,w_{i-p+1,1},w_{i,2},\ldots,w_{i-p+1,2},\ldots,w_{i-p+1,d})^{\top}$. Then Algorithm 1 can still be applied. Since the multivariate information is summarized by the distance matrix $\mathbf{L}=(l_{ij})_{0\leq i\leq n_{1},n_{1}+1\leq j\leq n-1}$, where $l_{ij}=L(\mathbf{v}_{i}(\mathbf{W}),\mathbf{v}_{j}(\mathbf{W}))$, regardless of the dimension $d$, step (a) in Algorithm 1 is identical as the scalar case. In the notable special case where $L(\mathbf{u},\mathbf{v})=\|\mathbf{u}-\mathbf{v}\|^{2}$, (5) becomes
@@
where $\|\cdot\|_{F}$ is the Frobenius norm. Straightforward calculations show that fixing $\mathbf{\Pi}=\hat{\mathbf{\Pi}}$, if $\hat{\mathbf{W}}$ is a solution to step (b) of Algorithm 1, then for each $l\in\{1,2,\ldots,d\}$, $(\hat{w}_{0,l},\hat{w}_{1,l},\ldots,\hat{w}_{n-1,l})$ also satisfies the system of linear equations (7) in the case of no side information. Therefore, in step (b) of Algorithm 1, each component in the multivariate time series can be imputed in parallel to lower the computational overhead and only the values of individual series are involved in computing the imputation in each component. The cross-sectional information is exploited when learning the optimal transport coupling in step (a). Thus, the algorithm is fast to implement in the multivariate setting.
@@
We shall compare TWI with the approach proposed by Carrizosa et al., (2013), which also aims to alleviate the corrupted statistical patterns caused by imputation. Their method is based on penalizing the deviation of some moments from the theoretical values. As a result, the moments, such as the autocorrelation functions, of the imputed series are close to the pre-specified values. However, this approach has a number of drawbacks. First, good theoretical values for the moments may not be readily available in practice, especially when the missing values are many. Second, their approach promotes smoothness of the time series and hence positive autocorrelation, which does not necessarily align with the underlying time series. Finally, the optimization problem is highly non-convex, leading to an intense computational hurdle. In contrast, the proposed TWI is data-driven, completely nonparametric, and enjoys the computational advantages brought by the alternating minimization algorithm and recent advances in computational optimal transport.
@@
Next, we discuss the imputation of nonstationary time series. For an $I(1)$ process $\{x_{t}\}$, $y_{t}=x_{t}-x_{t-1}$ is stationary and one can apply imputation methods to the first-differenced series $\{y_{t}\}$. More generally, one can apply TWI to the series $y_{t}=\phi(B)x_{t}$, where $\phi(z)=(1-z)^{d}$ if $x_{t}$ is an $I(d)$ process, and $B$ is the back-shift operator. Then, an imputation for $x_{t}$ can be recovered by inverting $\phi(B)$. Note that, in this case, the number of missing data in $\{y_{t}\}$ multiplies, and we need at least $d+1$ contiguous observed values in $\{x_{t}\}$ to obtain an observed value for $\{y_{t}\}$. To mitigate the information loss, we can encode the information from the raw data via the admissible set $\mathcal{C}$. For example, consider the $I(1)$ case: If $x_{t}$ is observed, then since $y_{1}+\ldots+y_{t}=x_{t}-x_{0}$, we may restrict the imputation $\{w_{t}\}$ for $\{y_{t}\}$ to satisfy $w_{1}+\ldots+w_{t}=x_{t}-x_{0}$ (assuming $x_{0}$ is observed). Finally, although we have motivated TWI from the stationarity of the underlying process, TWI fares quite well with certain nonstationary processes such as cyclic time series. In Section 5, we further explore the application of TWI to nonstationary time series.
@@ 4 Theoretical properties
In this section, we investigate the optimization and statistical properties of TWI. We first show Algorithm 1 converges to critical points. Then, asymptotic consistency of TWI is established when the missing data concentrate on one side of $n_{1}$. For general non-systematic missing patterns, we show that identification of the underlying marginal distributions can be achieved by TWI under the assumptions of interpolator stability using a two-state Markov chain example (see (C1) and (C2) in Section 4.2).
Before presenting the results, it is worthwhile to clarify the asymptotic regime under consideration. The analysis is meaningful only when the number of missing entries is non-negligible compared to the sample size. Otherwise, if $\sharp(\mathcal{M})=o(n)$, where $\sharp(E)$ is the cardinality of the set $E$, then the empirical marginal distributions of any (bounded) imputations satisfy
@@
where $\mathcal{A}_{n}=\{t\in\{p-1,\ldots,n-1\}:\{t,t-1,\ldots,t-p+1\}\cap\mathcal{M}\neq\emptyset\}$. Therefore, assuming $\sharp(\mathcal{M})=o(n)$ would essentially preclude distributional inconsistency in the downstream. To avoid this trivial case, in this paper we maintain instead the assumption that $\mathcal{M}=\mathcal{M}_{n}$ is a sequence of index sets (independent of the underlying time series) such that $\sharp(\mathcal{M}_{n})/n\asymp 1$, where $a_{n}\asymp b_{n}$ means there exists $0<\underline{c}<\bar{c}<\infty$ such that $\underline{c}\leq a_{n}/b_{n}\leq\bar{c}$ for all large $n$.
@@ 4.1 Optimization and statistical guarantees
We begin by studying the optimization properties of Algorithm 1. All proofs are relegated to Section S1 of the supplementary material.
@@
Suppose $\mathcal{C}$ is a convex set and $F(.,.)$ is defined in (5). Then,
@@
    (monotonicity) $F(\hat{\mathbf{w}}^{(t)},\hat{\mathbf{\Pi}}^{(t+1)})\leq F(\hat{\mathbf{w}}^{(t)},\hat{\mathbf{\Pi}}^{(t)})\leq F(\hat{\mathbf{w}}^{(t-1)},\hat{\mathbf{\Pi}}^{(t)})$, $t=1,2,\ldots$;
@@
    every limit point $(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})$ of $\{(\hat{\mathbf{w}}^{(t+1)},\hat{\mathbf{\Pi}}^{(t)}):t=1,2,\ldots\}$ is a critical point of $F$. i.e.,
@@
Due to the biconvexity of $F$, one can employ off-the-shelf solvers for the subproblems in Algorithm 1. Moreover, it follows directly from the proof that if $\lambda>0$, $F(\mathbf{w},\mathbf{\Pi})$ is $\lambda$-strongly convex in $\mathbf{w}$. Hence, step (b) in Algorithm 1 admits a unique solution in each iteration and can be solved efficiently.
A few remarks are in order. First, it is straightforward to generalize Proposition 3 for multivariate time series. Second, like most methods involving non-convex optimization, a good initialization is crucial for converging to a “good” critical point. In our simulation studies, we use simple linear interpolation or ARIMA-based Kalman smoothing as initializations and TWI is able to find an imputation that offers substantial improvements over these initializations. Third, unlike many non-convex optimization problems, we are not satisfied by an arbitrary global minimizer of $F$. Essentially, it is because among all processes that have stable $p$-dimensional marginal distributions, only a subset of them matches the underlying marginal distributions. In Section 4.2, we probe this identification issue in more detail. However, when the missing data happen to lie on the same side of $t=n_{1}$, the global minimizer of $F$ achieves asymptotic consistency in the sense of Proposition 4 below.
@@
Assume (A1) and (A2) in Section S1 of the supplementary material. Suppose for any $\mathbf{w}\in\mathcal{C}$, $w_{t}=\tilde{x}_{t}$ if $t\notin\mathcal{M}_{n}$. Suppose further $n$ and $n_{1}$ satisfy $n_{1}\rightarrow\infty$ and $n-n_{1}\rightarrow\infty$. If either $\mathcal{M}_{n}\subseteq\{0,1,\ldots,n_{1}\}$ or $\mathcal{M}_{n}\subseteq\{n_{1}+1,\ldots,n-1\}$ (but not both), then, the global minimizer $(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})$ of $F$ with $\lambda=0$ and $L(\mathbf{u},\mathbf{v})=\|\mathbf{u}-\mathbf{v}\|^{k}$, subject to $\mathbf{w}\in\mathcal{C}$ and the coupling conditions, satisfies
$\mathcal{W}_{k}\left(\frac{1}{n_{1}-p+2}\sum_{t=p-1}^{n_{1}}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})},\mu_{p}\right)+\mathcal{W}_{k}\left(\frac{1}{n-n_{1}-1}\sum_{t=n_{1}+1}^{n-1}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})},\mu_{p}\right)\rightarrow 0\quad\mbox{almost surely}.$